OpenStudy (kpop4life123):

The table shows the number of students in a school who like rap and/or hip-hop: Like Rap Do Not Like Rap Total Like Hip-Hop 35 8 43 Do Not Like Hip-Hop 10 21 31 Total 45 29 74

OpenStudy (kpop4life123):

What is the relative frequency, by row, of students who do not like rap or hip-hop? (4 points) 0.17 0.68 0.72 0.81

OpenStudy (kpop4life123):

@Addicted2Anime

OpenStudy (addicted2anime):

not sure on this one sorry

jhonyy9 (jhonyy9):

74 ----- 100% 39 ----- x ----------------- x = (100*39)/74 = 3900/74 = total there are 74 students - yes ? from these 74 are 8 do not like rap and 10+21 do not like hip-hop - total 31+8=39 there are 39 students who do not like rap or hip-hop - yes ? so than i think that we need calculi 39 is how many percent of 74 but i not have got a result from these choices - why ?